OpenStudy (anonymous):

-4x-2y=-12 , 4x+8y=-24 How do you solve this?

OpenStudy (anonymous):

what are you solving for?

OpenStudy (anonymous):

solving systems of equation by elimination

OpenStudy (anonymous):

so you arent solving for x or y?

OpenStudy (anonymous):

Yea I Am

OpenStudy (anonymous):

for both x and y?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

ok. to get x, you have to get the y on the other side. do you know how to do that?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

ok. well you have to do the opposite. since you are subtracting 2y on the left side, you add 2y on the right side. so the problem would be: -4x=-12+2y

OpenStudy (anonymous):

Never Mind g2g ...Thank you for the help though.

OpenStudy (anonymous):

ok no prob

OpenStudy (anonymous):

; )

OpenStudy (anonymous):

-4x - 2y = -12 4x + 8y = -24 ------------- 0 + 6y = - 36 6y = - 36 y = -36/6 y = - 6 now sub -6 in for y in either of the original equations 4x + 8y = - 24 4x + 8(-6) = - 24 4x - 48 = - 24 4x = -24 + 48 4x = 24 x = 24/4 x = 6 now lets check it... -4x - 2y = - 12 -4(6) - 2(-6) = - 12 - 24 + 12 = - 12 -12 = -12 (correct) ANSWER: x = 6 and y = -6 or (6,-6)